# S4ST: A Strong, Self-transferable, faSt, and Simple Scale Transformation for Data-free Transferable Targeted Attack

### v2

@@ Abstract
Transferable Targeted Attacks (TTAs), which aim to deceive black-box models into predicting specific erroneous labels, face significant challenges due to severe overfitting to surrogate models. Although modifying image features to generate robust semantic patterns of the target class is a promising approach, existing methods heavily rely on large-scale additional data and prolonged training, or at least require query access to victim models for method design and tuning. This dependence on prior information undermines the fair evaluation of TTA threats, potentially leading to a false sense of security or unnecessary overreactions. In this paper, we introduce two novel blind estimation measures—surrogate self-alignment and self-transferability—to analyze the effectiveness and correlations of basic transformations, to enhance gradient-based data-free attacks under strict black-box constraints (no access to victim models or their data distributions). Our findings challenge conventional assumptions: (1) Attacking simple scaling transformations uniquely enhances targeted transferability, outperforming other basic transformations and rivaling leading complex methods; (2) Geometric and color transformations exhibit high internal redundancy despite weak inter-category correlations. These insights drive the design and tuning of S4ST (Strong, Self-transferable, faSt, Simple Scale Transformation), which integrates dimensionally consistent scaling, complementary low-redundancy transformations, and block-wise operations. Extensive experiments on the ImageNet-Compatible dataset demonstrate that S4ST achieves a 77.7% average targeted success rate (tSuc) in the 10-Targets (all-source) setting, surpassing existing transformations (+17.2% over H-Aug with only 26% computational time) and state-of-the-art TTA solutions (+6.2% over SASD-WS with 1.2M samples for post-training). Notably, it attains 69.6% and 55.3% average tSuc against three commercial APIs and vision-language models, respectively. This work establishes a new SOTA for TTAs, highlights their potential threats, and calls for a reevaluation of the data dependency in achieving targeted transferability. Codes are available at https://github.com/scenarri/S4ST.
@@ 1 Introduction
Deep Neural Networks (DNNs) have been shown to be vulnerable to adversarial examples (AEs) [1], which refer to images with small, intentionally crafted perturbations that are usually imperceptible to the human visual system but can fool DNNs. The presence of AEs raises significant concerns when deploying DNNs in security-critical domains, including autonomous driving, face recognition, and medical image analysis. Therefore, finding, understanding, and defending AEs have received significant attention [2, 3, 4, 5]. AEs exhibit an intriguing property known as transferability [6], which means that AEs designed for one DNN can successfully deceive other DNNs with distinct architectures. The transferability of AEs facilitates the implementation of black-box attacks where queries to the victims are forbidden. As a result, significant attention has been paid to enhancing these examples’ transferability in the last decade [7].
While transferable untargeted attacks have been extensively investigated, yielding a substantial body of research [8, 9, 10, 11, 12, 13], Targeted Transferable Attacks (TTAs) are significantly more challenging and consequently less explored. This is primarily due to the exacerbated overfitting of AEs to surrogate models in the context of targeted attacks. More specifically, it is challenging to construct a targeted perturbation that can drive samples to a specific target class within the decision spaces of various unknown models. To achieve effective targeted transferability, directly modifying the original visual features to form robust semantic patterns of the target class has become a mainstream alternative [14, 15], requiring the exploitation of extensive data to model the underlying distribution. To this end, data-reliant TTAs often entail optimizing universal perturbations [16], generators [15, 17, 18], or the surrogate model itself [19, 20, 21]. However, two key drawbacks exist due to their need for tens of thousands of additional data points:
@@
    The data used to model the distribution is typically the training or testing data of the target model, leading to data leakage and violating the black-box attack setting. This hinders the proper evaluation of threats by the research community.
@@
    Most of these methods demand substantial time and resources for training. For instance, generation-based methods require training a separate generator for each target class, which limits the further utilization of targeted adversarial examples such as cross-dataset applications and adversarial training.
In this paper, we focus on data-free TTAs, also known as the gradient-based attacks [8, 9], which involve executing gradient descent on any (sample, label) pair, offering both efficiency and flexibility. In the meantime, these are plagued by severe overfitting and poor transferability [22].
The limitations of data-free TTAs have been primarily attributed to the gradient vanishing issue of the cross-entropy loss function in mainstream research, which has spurred the development of a series of alternative loss functions [23, 24, 25, 26]. However, input transformations play a more critical role [10, 27]. These transformations are incorporated into the iterative gradient update process, serving a dual purpose: on one hand, they diversify the image features, and on the other hand, they soften the surrogate’s prediction distribution, thereby mitigating gradient vanishing. This dual mechanism makes the realization of effective data-free TTAs possible. In fact, specialized loss functions lose their efficacy when decoupled from input transformations [24]. Some existing transformations designed for untargeted attacks such as the random scaling & padding transformations in [10] and [28], have demonstrated promising improvements [29, 24]. Nevertheless, there is still a lack of guidelines for designing transformation methods under black-box conditions, and their potential remains largely locked. Therefore, we aim to understand how to design highly effective and efficient transformation methods, to what extent these methods can unlock the potential of data-free TTAs, and whether it is possible to outperform data-reliant methods without access to the target model and its training data.
In this paper, we simplify the above objective to identify a strategy to combine various basic transformations. First, we propose a clever use of the overfitting of targeted AEs. Specifically, the more severe the overfitting, the weaker the impact of AEs on unknown models, indicating a greater misalignment between the surrogate model and unknown models on AEs. Therefore, we directly use the self-alignment of the surrogate model on transformed AEs and clean samples to estimate the alignment towards unknown models on AEs, i.e., the effectiveness of basic transformations. Furthermore, by attacking the surrogate model using various complex transformation methods and computing its self-transferability against basic transformation defenses, we explore the correlations (or redundancies) among basic transformations. Through these two simple yet effective analyses, the following key findings are gained:
@@
    Simple scaling transformations exhibit outstanding capability in promoting targeted transferability. Compared to other basic transformations, scaling demonstrates a stronger effect on promoting surrogate self-alignment and operates effectively over a wider range of transformation intensities.
@@
    Geometric and color transformations exhibit high internal similarity but weak inter-category correlation. That is, even if two transformations appear to be less related, there still exists strong redundancy, such as between scaling and rotation transformations. This observation reminds us to cautiously seek complementary enhancements.
Figure 1: Comparison against existing transformation methods at incremental attack iterations. Compared with the previous SOTA transformation, H-Aug [27], S4ST yields an absolute improvement of 14.2% on average tSuc to 83.0%, and a relative reduction of 74.3% on time consumption to 1.45s. Additionally, compared with the previous setting ($T=300$), more attack iterations benefit a lot for potent transformations like our S4ST.
Further, the findings above have enabled the development of a novel and powerful scaling-centered transformation method for data-free TTAs, named Strong, Self-transferable, faSt, and Simple Scale Transformation (S4ST). S4ST amalgamates three efficient strategies to enhance the simple scaling transformation: 1) Modifications that maintain dimensional consistency while injecting greater diversity and reducing computational burden; 2) Integration with complementary transformations that are less correlated with scaling; and 3) Implementation of block-wise scaling. It is worth noting that the proposed S4ST requires no extra data to determine what complementary transformations should be applied and, even for parameter tuning (all done according to the self-transferability analysis in a strict black-box manner)!
In summary, we provide novel insights and methods for designing input transformation methods to enhance TTAs, and we show that under strict black-box conditions—where any access of the black-box model or its data is prohibited—it is possible to complete method design and parameter tuning while achieving SOTA TTA performance. Our work highlights the importance of dedicating resources to the evaluation and defense against targeted AEs in safety-critical applications. Overall, the main contributions of this paper are as follows:
@@
    We propose two blind estimation measures: surrogate self-alignment and self-transferability against basic transformations. Serving as effective proxies for the black-box transferability of targeted AEs, they enable feasible analyses of the effectiveness and synergistic effect of various basic transformations without accessing any additional data and victim models, significantly diminishing the dependence on empirical or intuitive choices common in the prior art.
@@
    To the best of our knowledge, we are the pioneers in revealing the unique superior efficacy of the simple scaling transformation in enhancing targeted transferability, and the redundancies within geometric and color transformations. Based on these, we further design S4ST, an advanced scaling-centered transformation that ingeniously integrates modified scaling with complementary transformations and leverages the benefits of block-wise operations, adhering to a strict black-box manner.
@@
    Extensive and comprehensive experiments on the ImageNet-Compatible dataset substantiate the effectiveness and efficiency of S4ST. It not only outperforms existing transformation techniques in terms of both effectiveness and efficiency (see Fig. 1) but also surpasses SOTA TTA solutions that depend on 50k to 1.2 million training samples (see Table VII), all without requiring additional data. It also exhibits exceptional transferability to commercial APIs and vision-language models. It is found to significantly improve the exploitation on existing visual elements to produce highly transferable target semantics, circumventing the necessity to learn this capability from extensive multi-class data.
The rest of this paper is organized as follows. Section 2 introduces the problem definition, related works, and experimental validations of existing consensuses in designing transformation methods. Section 3 elaborates on the motivation, analysis, findings, and details of our method. Section 4 offers the experimental settings, results, and analysis. Finally, Section 5 concludes this work.
@@ 2.1 Problem Setting and Baseline Approach
Consider an image classifier $F$ that predicts a label among $K$ categories, $\bm{y}=\{y_{i}|y_{i}\in\mathbb{Z},0\leq i<K\}$, for an input with RGB channels, $\bm{x}\in\mathbb{R}^{3\times H\times W}$, where $H$ and $W$ denote the image height and width, respectively. Without loss of generality, we denote $F=C\circ f$, where $f(\bm{x})=\bm{p}\in[0,1]^{K}$ represents the prediction distribution and $C(\bm{p})=y_{\operatorname{argmax}(\bm{p})}$ indexes the label with the highest probability. TTA aims to manipulate black-box classifiers, $G=C\circ g$, to produce an arbitrary targeted prediction, $y_{t}$, for any input:
@@
Here, The step size $\alpha$ and perturbation budget $\epsilon$ are configured to 2/255 and 16/255, respectively. The momentum accumulation decay factor $\mu$ is set to 1 [8], and a 5$\times$5 Gaussian kernel [9] $\bm{W}$ is used for TI. We adopt the margin-based cross-entropy loss [25] and the attack iteration is extended to $T=900$ (compared to 300 in previous work) in keeping with the increased diversity afforded by the most advanced transformation methods $\mathcal{T}$.
@@ 2.2 Transferable Targeted Attacks
In this paper, current solutions for TTA are categorized into data-free and -reliant approaches as elaborated upon in the following.
@@ 2.2.1 Data-free TTAs
Data-free TTAs employ gradient-based iterative optimization on a target (sample, label) pair using a surrogate model [30, 8, 9, 2]. While effective in untargeted attacks, the cross-entropy loss faces challenges in TTAs due to gradient vanishing issues [23, 24]. Let $\bm{\phi}$ denote the representation produced by $f$, such as features preceding the final linear layer. The logit values are computed as $l_{i}=w_{i}^{\top}\bm{\phi}+b_{i}$, with $w_{i}$ and $b_{i}$ being the class-specific weight and bias, respectively. The normalized probability with respect to each category is given by the SoftMax function: $p_{i}=\frac{e^{l_{i}}}{\sum e^{l_{j}}}$. The gradient
Beyond the design of loss functions, it is the image transformations, as elaborated upon below and in Sec. 2.3, that have proven crucial for achieving effective data-free TTAs and have so far been underexplored. Zhao et al. [24] demonstrated that the Diverse Inputs (DI) transformation [10] significantly enhances the transferability of targeted attacks, demonstrating that AEs rarely transfer successfully without it. An exception can be found in [29] where targeted transferability is boosted via mixing the clean and adversarial features at intermediate DNN layers. However, it also largely depends on transformation techniques. Although we categorize the aforementioned methods as data-free, it is important to emphasize that they still require leveraging feedback from black-box models to tune their parameters. The optimization process was observed to require substantially more iterations (approximately 300) to converge. Since then, the DI method has become a standard benchmark in subsequent research. However, efforts to understand which transformations are essential for targeted transferability remain scarce. Wei et al. [27] provided empirical evidence suggesting that DI-transformed images diversify the discrimination regions of the surrogate model across iterations, thereby expanding the final discrimination region with respect to the targeted label in black-box models. They also introduced the T/H-Aug method, which applies random combinations [31] of attention-deviation transformations. Byun et al. proposed the Object-based DI (ODI) [32] transformation, inspired by the intuition that AEs transferable across various 3D shapes and viewpoints could more effectively compromise black-box models. In [33], the authors add an extra inference branch to pursue a consistent adversarial effect between random crops and the whole image.
In this work, we explore in greater depth the role of image transformations within data-free TTAs. We also challenge the limitations of existing empirical insights and intuitive approaches, proposing new methods to analyze various basic transformations. Building on these, we have developed the S4ST transformation, which has led to notable performance improvements in a strict black-box setting.
@@ 2.2.2 Data-reliant TTAs
Data-reliant TTAs seek to improve targeted transferability by leveraging the data distribution through a substantial volume of additional data. Targeted universal attacks [14, 16] develop a single perturbation across a large dataset, enabling the universal attack of multiple unseen samples and demonstrating cross-model transferability. Inkawhich et al. chose to train auxiliary models at intermediate layers to capture and attack the feature distribution relative to the target class [34]. Springer et al. reveal that adversarially training surrogate model against small-magnitude AEs yields substantial improvement in targeted transferability [35]. The score-matching techniques are borrowed by Intrinsic Adversarial Attack (IAA) [19] and Distribution-Relevant Attack (DRA) [20] to respectively tune the surrogate models’ architectural parameters and weights for modeling the data distribution. Wu et al. propose the Sharpness-Aware Self-Distillation and Weight Scaling (SASD-WS) method, which combines the sharpness-aware optimization [36] and model distillation for model self-enhancement and results in better transferability.
The pioneering generative framework for TTA, known as Transferable Targeted Perturbation (TTP) [15], involves training a generator on extensive data to align with a specified target class in terms of both global distribution and local neighborhood structure. Zhao et al. introduced the Minimizing Maximum Model Discrepancy (M3D) attack [18], which discriminates and targets two surrogate models (the original and its duplicate). Conditional Generative Semantic Pattern (C-GSP) [37] enables perturbation generators to attack multiple labels simultaneously by incorporating conditional learning.
Although these methods are effective and efficient in inference, their practicality is constrained by the requirement for extensive training data from the same domain or even the same dataset as the black-box models and the laborious retraining process for large amounts of target labels. By building on our proposed S4ST transformation, data-free TTAs have surprisingly surpassed SOTA data-reliant techniques, all while preserving remarkable flexibility and effectiveness.
2.3 Input Transformations for Transferability
Image transformation is an effective and necessary strategy in training deep vision models, effectively mitigating the need for extremely large quantities of data and annotations [38, 39, 40]. For example, horizontal flips and random resized cropping are commonly used in natural image understanding tasks, significantly boosting generalization performance. Furthermore, transformation strategies are typically data-dependent and can be conveniently implemented in a plug-and-play manner across different model architectures. The same benefits extend to the domain of adversarial attacks. There are two primary streams of transformation strategies for enhancing adversarial transferability: 1) Dynamically searching for optimal combinations from a vast pool of transformations during the attack iteration [41, 42, 43]; and 2) Consistently applying a predefined transformation in each iteration. In this paper, we focus exclusively on the latter to avoid the need to access black-box models for optimization. In addition to TTA-oriented transformation methods [27, 32], we also examine several prominent works in the field of untargeted attacks. This is to validate whether the prevalent consensuses in transformation design are applicable in the context of TTA. An overview of the methods involved is presented in Fig. 2, introduced below.

### v3

@@ Abstract
Transferable Targeted Attacks (TTAs) face significant challenges due to severe overfitting to surrogate models. Recent breakthroughs heavily rely on large-scale training data of victim models, while data-free solutions, i.e., image transformation-involved gradient optimization, often depend on black-box feedback for method design and tuning. These dependencies violate black-box transfer settings and compromise threat evaluation fairness. In this paper, we propose two blind estimation measures, self-alignment and self-transferability, to analyze per-transformation effectiveness and cross-transformation correlations under strict black-box constraints. Our findings challenge conventional assumptions: (1) Attacking simple scaling transformations uniquely enhances targeted transferability, outperforming other basic transformations and rivaling leading complex methods; (2) Geometric and color transformations exhibit high internal redundancy despite weak inter-category correlations. These insights drive the design and tuning of S4ST (Strong, Self-transferable, faSt, Simple Scale Transformation), which integrates dimensionally consistent scaling, complementary low-redundancy transformations, and block-wise operations. Extensive evaluations across diverse architectures, training distributions, and tasks show that S4ST achieves state-of-the-art effectiveness-efficiency balance without data dependency. We reveal that scaling’s effectiveness stems from visual data’s multi-scale nature and ubiquitous scale augmentation during training, rendering such augmentation a double-edged sword. Further validations on medical imaging and face verification confirm the framework’s strong generalization. Codes are available at https://github.com/scenarri/S4ST.
@@ 1 Introduction
Deep Neural Networks (DNNs) have been shown to be vulnerable to adversarial examples (AEs) [67], which refer to images with small, intentionally crafted perturbations that are usually imperceptible to the human visual system but can fool DNNs. The presence of AEs raises significant concerns when deploying DNNs in security-critical domains, including autonomous driving [75, 5], face recognition [2], and remote sensing [54], regarding both digital and real-world [62] scenarios. Therefore, finding, understanding, and defending AEs have received significant attention [48, 73, 93, 34]. Crucially, AEs exhibit transferability [21], successfully fooling unseen victims based on a surrogate model. This property enables black-box transfer attacks, driving efforts to enhance transferability [22].
While untargeted transferability is well-studied, yielding a substantial body of research [16, 17, 90, 78, 76, 41], Targeted Transferable Attacks (TTAs) are significantly more challenging and consequently less explored. This is primarily due to the difficulties of driving samples into specific target decision spaces across unknown architectures. To form robust target semantics [96, 50], representative methods optimize universal perturbations [85], generators [50, 99], or the surrogate model itself [104, 103, 87]. However, relying on extensive data to model distributions categorizes them as data-reliant TTAs, presenting two major drawbacks:
@@
    They often exploit victim training/testing data, causing data leakage that violates strict black-box settings and impedes fair threat evaluation.
@@
    They require massive training resources. For instance, generative attacks [99] train class-specific generators, hindering broader applications like cross-dataset transfer or adversarial training.
In this paper, we focus on data-free TTAs, also known as the gradient-based attacks [16, 17], which efficiently optimize any (sample, label) pair but suffer from severe overfitting and poor transferability [44]. While often attributed to cross-entropy gradient vanishing, prompting alternative loss functions [37, 100, 84, 64], input transformations actually play a more critical role [90, 83]. As Zhao et al. [100] showed, AEs rarely transfer without them. Crucially, despite being termed “data-free”, existing transformation methods still rely on heuristic choice or implicit black-box feedback for design and tuning. Guidelines for strict black-box transformation design remain absent, leaving their potential largely untapped.
Figure 1: What drives the method design and training distinguishes this study from existing work: (a) Data-reliant TTAs, e.g., generative [50, 99] and universal [85] attacks require large-scale data (often the very training data of the black-box models) and extensive training for each target label; (b) Data-free TTAs, e.g., transformation-based gradient attacks, typically depend on query access or feedback from the black-box model for method design and tuning; (c) Our solution introduces two estimation metrics, self-alignment and self-transferability, to analyze the effectiveness and correlations of basic transformations. This enables the design of composite transformations and facilitates effective attacks without any additional data or access to the black-box model.
Figure 2: Conceptual illustration of the proposed self-alignment (left) and self-transferability (right) blind estimation measures.
This work explores TTAs under strict black-box constraints using only a domain-matched surrogate model [17, 16, 50, 99]. Unlike existing solutions (Fig. 1), we aim to design data-free transformations to prevent data leakage and ensure fair threat assessment. This requires integrating basic transformations effectively by analyzing per-transformation effectiveness and cross-transformation correlations. As conceptually illustrated in Fig. 2, since overfitting bottlenecks transferability, we first leverage the surrogate’s self-alignment to estimate a transformation’s ability to restore the representation distribution, thereby preserving transferability toward unknown black-box models. Intuitively, this can be viewed in an adversarial learning fashion: we seek the most challenging virtual victim, the surrogate prepended with a transformation, $\bm{\phi}\circ\mathcal{T}$, for the AE optimization. Furthermore, we evaluate cross-transformation correlations via self-transferability. Attacking the surrogate with complex transformations and evaluating against basic transformations (virtual defenses) allows us to objectively identify synergies and eliminate redundancy. These surrogate-internal measures reveal key findings:
@@
    Simple scaling outstandingly promotes targeted transferability. It achieves stronger surrogate self-alignment and operates effectively across broader intensity ranges than other basic transformations.
@@
    Geometric and color transformations show high internal similarity but weak inter-category correlation. Strong redundancy exists even between seemingly unrelated transformations (e.g., scaling and rotation), necessitating cautious selection for complementary enhancements.
These findings drive the development of S4ST (Strong, Self-transferable, faSt, and Simple Scale Transformation), a novel scaling-centered transformation for data-free TTA. S4ST enhances simple scaling via three efficient strategies: 1) dimensionally consistent modifications injecting diversity while reducing computation; 2) integration with low-correlation complementary transformations; and 3) block-wise scaling. Notably, S4ST determines complementary transformations and tunes parameters entirely via self-transferability analysis in a strict black-box manner, requiring no extra data while achieving a superior effectiveness-efficiency balance (Fig. 3).
We further discover that scaling’s distinct effectiveness roots in visual data’s multi-scale nature and modern deep vision practices. Standard training universally employs scale augmentations (e.g., RandomResizedCrop), embedding scale invariance as a shared inductive bias across most vision models. S4ST strategically exploits this to craft highly transferable AEs without violating black-box constraints.
Figure 3: Comparison against existing transformation methods at incremental attack iterations (ImageNet-Compatible dataset). Compared with the previous SoTA, H-Aug [83], S4ST yields an absolute improvement of 14.2% on average tSuc to 83.0%, and a relative reduction of 74.3% on time consumption to 1.45s. Compared with the previous setting ($T=300$), more attack iterations benefit a lot for potent transformations like our S4ST. tSuc denotes targeted attack success rate and experimental details are in Section 4.3.1.
In summary, we demonstrate that under strict black-box conditions, prohibiting any access to victim models and their training data, it is possible to design and tune transformation methods that achieve state-of-the-art (SoTA) targeted transferability. Our main contributions are:
@@
    We propose self-alignment and self-transferability as blind estimation measures. They serve as effective proxies to analyze basic transformations’ effectiveness and synergies without accessing victims or extra data, reducing the reliance on empirical choices common in prior art.
@@
    We pioneer the discovery of simple scaling’s superior efficacy in enhancing targeted transferability. This stems from visual data’s inherent nature and the universal adoption of scale augmentation during training, revealing a dual-edged sword: practices enhancing generalization simultaneously introduce transfer attack vulnerabilities.
@@
    We propose S4ST, an advanced scaling-centered transformation integrating modified scaling, complementary transformations, and block-wise operations under strict black-box constraints.
  4. 4.
    Extensive evaluations across natural images, medical imaging, and face verification validate our framework’s transferability. S4ST outperforms existing transformation methods and data-reliant SoTA TTAs (using 50k-1.2M samples), showing robust transferability to commercial APIs and vision-language models (VLMs).
Aligning with advancements in adversarial machine learning, this research aims to deepen our understanding of adversarial vulnerabilities, thereby contributing to the development of trustworthy artificial intelligence systems [52, 91]. The concept of targeted transferability is not only instrumental in enhancing privacy and intellectual property protection [19, 10] and enabling robust inference [48] but also critical for identifying potential risks in high-stakes applications such as autonomous driving and VLMs [43]. Although malicious exploitation is possible, we hope this work catalyzes the development of robust defenses and positive societal impacts.
The remainder is organized as follows: Section 2 introduces the problem, related works, and existing consensus validations. Section 3 details our motivation, analysis, and methodology. Sections 4 and 5 present experiments and discussions. Section 6 concludes the paper.
@@ 2.1 Problem Setting and Baseline Approach
Consider an image classifier $F$ that predicts a label among $K$ categories, $\bm{y}=\{y_{i}|y_{i}\in\mathbb{Z},0\leq i<K\}$, for an input with RGB channels, $\bm{x}\in\mathbb{R}^{3\times H\times W}$, where $H$ and $W$ denote the image height and width, respectively. Without loss of generality, we denote $F=C\circ\bm{\phi}$, where $\bm{\phi}(\bm{x})\in\mathbb{R}^{K}$ represents the logits and $C(\cdot)$ indexes the label with the highest probability. $f(\bm{x})=\operatorname{Softmax}(\bm{\phi}(\bm{x}))=\bm{p}_{\bm{\phi}}\in[0,1]^{K}$ represents the final predction distribution. TTA aims to manipulate black-box classifiers, $G=C\circ\bm{\psi},g(\bm{x})=\bm{p}_{\bm{\psi}}$, to produce an arbitrary targeted prediction, $y_{t}$, for any input:
@@
The step size $\alpha$ and perturbation budget $\epsilon$ are configured to 2/255 and 16/255, respectively. The accumulation decay factor $\mu$ is set to 1 for momentum $\bm{g}$ [16], and a $\times5$ Gaussian kernel [17] $\bm{W}$ is used for TI. We adopt the margin-based cross-entropy loss [84] and the attack iteration is extended to $T=900$ (compared to 300 in previous work) in keeping with the increased diversity afforded by the most advanced transformation methods $\mathcal{T}$.
@@ 2.2 Transferable Targeted Attacks
We categorize existing TTAs into data-free and data-reliant paradigms, analyzing their limitations to contextualize the distinct contributions of S4ST.
Figure 4: Average tSuc evaluated on the ImageNet-Compatible. Input transformations are identified as key role in enabling effective data-free TTAs.
@@ 2.2.1 Data-free TTAs
Data-free TTAs typically rely on gradient-based iterative attacks [35, 16, 17, 48] using a surrogate model. Overcoming gradient vanishing (see supplementary) is a key challenge, addressed by novel loss functions that replace (e.g., PoTrip [37], Logit [100]) or soften the standard cross-entropy [84, 64]. Crucially, image transformations are essential for data-free TTAs. As Fig. 4 shows, specialized losses fail without them. Although untargeted transformations like random scaling & padding in DI [90] and RDI [105] show promise [3, 100], understanding which transformations specifically drive targeted transferability remains scarce.
Wei et al. [83] showed that DI [90] diversifies surrogate discrimination regions across iterations, expanding them in black-box models, and proposed T/H-Aug using random combinations of attention-deviation transformations. Byun et al. introduced Object-based DI (ODI) [2], hypothesizing that AEs transferable across 3D shapes and views better compromise black-box models.
However, these techniques rely on empirical observations or victim access, lacking a principled black-box understanding. To address this, our framework introduces self-alignment and self-transferability metrics to systematically evaluate transformations, challenge existing consensuses (Section 2.3), and analytically guide S4ST’s design. We also include prominent untargeted attack transformations (Fig. 5, details in supplementary) to validate whether prevalent design consensuses apply to TTAs.
@@ 2.2.2 Data-reliant TTAs
Data-reliant TTAs leverage large-scale data to enhance transferability via three main paradigms: 1) Targeted universal attacks [96, 85] optimize a single perturbation across large datasets to universally deceive models on unseen samples; 2) Surrogate refinement [63, 87, 104, 103] fine-tunes surrogate models for improved smoothness or data distribution matching; and 3) Generative attacks [50, 99, 92] train generators to model data distributions, facilitating AE transfer.
Despite their efficacy, these methods incur prohibitive data and computational costs. We demonstrate that such dependencies are unnecessary. For instance, while the generative M3D [99] requires $\sim$10 hours and 50K samples to train a per-label generator (see supplementary), our data-free S4ST achieves superior transferability with minimal overhead and ultimate flexibility for any (sample, label) pair.
2.3 Limitations of Existing Design Consensuses
We scrutinize the prevailing consensuses in developing transformations to assess their robustness and utility. We evaluate the black-box transferability of various transformations across default and expanded hyperparameter configurations. This transferability is measured by the target class ($t$) probability variance between adversarial (generated by $\mathcal{T}$-TMI) and clean samples:
We then examine correlations between $\textit{T}_{\mathcal{T}}$ and established consensus metrics: Diversity [77, 90, 40, 17, 78] (average cross-entropy loss on transformed inputs), Attention Deviation [17, 76, 83, 81, 39] (changes in model attention maps), and Gradient Magnitude [100, 82, 64, 84] (average $\ell_{2}$-norm of attack gradients). Details are in the supplementary material.